# Probing scalar non-standard interaction of supernova neutrinos in next-generation neutrino experiments

### v1

@@ 5 Summary and conclusion
This work investigates the flavor evolution of neutrinos in a core-collapse supernova, focusing on the effects of a scalar-mediated, flavor-conserving non-standard interaction between the outgoing neutrinos and the background matter particles alongside the standard model interactions. The new interactions introduce a correction term to the standard Dirac neutrino mass matrix, thereby modifying the neutrino propagation. By studying the evolution of the effective mixing matrix elements, we find that the presence of the SNSI interaction characterized by $\eta_{ee}$ does not have any notable impact on the neutrino flavor evolution, with $\nu_{e}$ alignment with effective mass state $\nu_{3,m}$ ($\nu_{2,m}$) within the supernova core in the NMO (IMO) scenario. In contrast, the presence of $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ leads to a qualitatively different behavior: due to a new level-crossing induced by the interaction, $\nu_{e}$ transitions to $\nu_{2,m}$ ($\nu_{1,m}$) for normal (inverted) mass ordering. We also find that the evolution of $\bar{\nu}_{e}$ remains largely unaffected by these new interactions.
We estimate the incoming $\nu_{e}$ and $\bar{\nu}_{e}$ fluxes at Earth from a core-collapse supernova originating from a $25M_{\odot}$ progenitor located at a distance of 10 kpc. We compute the estimated event distribution as a function of the time after the core bounce at the next-generation neutrino experiments DUNE and Hyper-K. At DUNE, $\nu_{e}$ events are primarily detected through charged-current (CC) interactions with argon nuclei, while Hyper-K is sensitive to $\bar{\nu}_{e}$ via inverse beta decay. Our results show that in DUNE, the expected event rate is significantly higher in the presence of SNSI parameter $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ during the initial time after the core bounce, for both the mass ordering of neutrino. However, a new degeneracy in the event distribution arises in the presence of the new interaction — the event distribution for normal mass ordering with SNSI matches that of the inverted mass ordering with standard interactions. We demonstrate that $\bar{\nu}_{e}$ events observed at the Hyper-K detector may help to lift the degeneracy. In the Hyper-K detector, we observe that for the IMO scenario, the expected event rates in each time bin are larger than those in the NMO scenario. However, the SNSI parameters do not have any impact on the $\bar{\nu}_{e}$ event distribution. Therefore, a combined analysis can help distinguish between the NMO+SNSI and the IMO+SI scenarios. We also discuss how event distribution in Hyper-K can help differentiate between some vector-mediated NSI and scalar-mediated NSI scenarios.
@@ Acknowledgments
We would like to thank Sudip Jana, Yago Porto, Bhaskar Dutta, Aparajitha Karthikeyan, and Nityasa Mishra for discussions. S.D. thanks the organizers of the PHENO-2025 conference at the University of Pittsburg for providing an opportunity to present the preliminary results of this work. This work is supported by the U.S. Department of Energy Grant DE-SC-0010113.

### v2

@@ 5 Summary and conclusion
This work investigates the flavor evolution of neutrinos in a core-collapse supernova, focusing on the effects of a scalar-mediated, flavor-conserving non-standard interaction between the outgoing neutrinos and the background matter particles alongside the standard model interactions. The new interactions introduce a correction term to the standard Dirac neutrino mass matrix, thereby modifying the neutrino propagation. By studying the evolution of the effective mixing matrix elements, we find that the presence of the SNSI interaction characterized by $\eta_{ee}$ does not have any notable impact on the neutrino flavor evolution, with $\nu_{e}$ alignment with effective mass state $\nu_{3,m}$ ($\nu_{2,m}$) within the supernova core in the NMO (IMO) scenario. In contrast, the presence of $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ leads to a qualitatively different behavior: due to a new level-crossing induced by the interaction, $\nu_{e}$ aligns with $\nu_{2,m}$ ($\nu_{1,m}$) for normal (inverted) mass ordering as long as $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ is greater than 0.01. We also find that the evolution of $\bar{\nu}_{e}$ remains largely unaffected by these new interactions.
We estimate the incoming $\nu_{e}$ and $\bar{\nu}_{e}$ fluxes at Earth from a core-collapse supernova originating from a $25M_{\odot}$ progenitor located at a distance of 10 kpc and 20 kpc. However, for 20 kpc scenario, expected event rate is four times smaller compared to 10 kpc case, as event rate is proportional to $1/R^{2}$, where $R$ is the distance between the Earth and the supernova. We compute the estimated event distribution as a function of the time after the core bounce at the next-generation neutrino experiments DUNE and Hyper-K. At DUNE, $\nu_{e}$ events are primarily detected through charged-current (CC) interactions with argon nuclei, while Hyper-K is sensitive to $\bar{\nu}_{e}$ via inverse beta decay. Our results show that in DUNE, the expected event rate for neutrinos is significantly higher in the presence of SNSI parameter $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ during the initial time after the core bounce, for both mass orderings of neutrinos. Considering 20% flux uncertainties and statistical uncertainties, we find that DUNE would be able distinguish between the SI and SNSI scenario if the mass ordering is known. However, a new degeneracy in the event distribution arises in the presence of a new scalar-mediated interaction — the event distribution for normal mass ordering with SNSI matches that of the inverted mass ordering with standard interactions. We demonstrate that $\bar{\nu}_{e}$ events observed at the Hyper-K detector may help to lift the degeneracy if the supernova distance is small and neutrino luminosity is large enough. In the Hyper-K detector, we observe that for the IMO scenario, the expected event rates in each time bin are larger than those in the NMO scenario. However, the SNSI parameters do not have any impact on the $\bar{\nu}_{e}$ event distribution. Therefore, a combined analysis can help distinguish between the NMO+SNSI and the IMO+SI scenarios.
@@ Acknowledgments
We would like to thank Sudip Jana, Yago Porto, Bhaskar Dutta, Aparajitha Karthikeyan, and Nityasa Mishra for discussions. We also thank Damiano F. G. Fiorillo and Edoardo Vitagliano for insightful comments. S.D. thanks the organizers of the PHENO-2025 conference at the University of Pittsburg for providing an opportunity to present the preliminary results of this work. This work is supported by the U.S. Department of Energy Grant DE-SC-0010113.
Appendix A Adiabatic propagation neutrinos in the presence of scalar NSI
For adiabatic neutrino propagation through a non-uniform medium like supernova, adiabaticity condition must be satisfied at the resonance. In this appendix, we test the adiabaticity condition for neutrino propagating from the supernova core to the surface in the presence of scalar NSI. Considering ($e,\mu$) block of a three-flavor neutrino Hamiltonian, the adiabaticity condition is given by Jana:2024lfm
$\gamma=\left|\frac{4|\mathcal{H}_{e\mu}|^{2}}{\dot{\mathcal{H}}_{ee}-\dot{\mathcal{H}}_{\mu\mu}}\right|>1\,$, (23)
where $\mathcal{H}_{\alpha\beta}$ denotes the elements of the Hamiltonian defined in eq. (6), which we further expand,
$\mathcal{H}=\frac{1}{2E}\left[MM^{\dagger}+\delta MM^{\dagger}+\delta M^{\dagger}M+\delta M\delta M^{\dagger}+2E\cdot V_{\rm SI}\right]\,$, (24)
where $M$, $\delta M$, and $V_{\rm SI}$ are defined in section 2. The first term in eq. (24) shows the vacuum part of Hamiltionian, whereas the other three terms are comprised of the standard matter interaction and the SNSI. In the inner region of the supernova, where the matter density is large, the matter interaction part of the Hamiltonian dominates. Considering only the SNSI parameter $\eta_{\mu\mu}$ to be non-zero, and neglecting the vacuum terms in the Hamiltonian, we get
$\mathcal{H}_{ee}$ $\simeq V_{\rm CC}=\sqrt{2}G_{F}N_{e}=\sqrt{2}G_{F}\frac{\rho}{m_{n}}Y_{e}\,$, (25)
$\mathcal{H}_{e\mu}$ $\simeq\frac{1}{2E}M_{e\mu}\delta M_{\mu\mu}=\frac{1}{2E}M_{e\mu}\delta\tilde{M}_{\mu\mu}\frac{\rho}{\rho_{s}}\,$, (26)
$\mathcal{H}_{\mu\mu}$ $\simeq\frac{1}{2E}\delta M_{\mu\mu}\delta M_{\mu\mu}=\frac{1}{2E}\left[\delta\tilde{M}^{2}_{\mu\mu}\left(\frac{\rho}{\rho_{s}}\right)^{2}\right]\,.$ (27)
$M_{\alpha\beta}$ ($\delta\tilde{M}_{\alpha\beta}$) denotes elements of the matrix $M$ ($\delta\tilde{M}$), defined in eq. 7. Note that in the expression for $\mathcal{H}_{\mu\mu}$, we neglect the term proportional to $\delta M$, which is small compared to the term with $\delta M_{\mu\mu}\delta M_{\mu\mu}$ at large densities. It is clear that the denominator in eq. (23) is proportional to $d\rho/dr$, considering $Y_{e}$ to be constant. Following the density profile in ref. Serpico:2011ir ; Jana:2024lfm , which approximately follows $\rho(r)\approx\rho_{c}\left(\frac{r}{R_{c}}\right)^{-3}$, where $\rho_{c}=10^{14}$ g/cc and $R_{c}=10$ km, we estimate $\gamma$ in eq. (23). After simplification, approximate expression for the $\gamma$ is,
$\gamma\simeq\frac{6.25\times 10^{-6}(\frac{r}{R_{c}})^{-6}}{\left|-3.2\times 10^{-10}\left(\frac{r}{R_{c}}\right)^{-4}+1.5\times 10^{-4}\left(\frac{r}{R_{c}}\right)^{-7}\right|}\,.$ (28)
We check that adiabaticity condition is satisfied at the density corresponding to the new resonance in the presence of SNSI parameters, ($\rho\approx 10^{8}~(10^{10})$ g/cc for $\eta_{\mu\mu}=0.1~(0.01)$) So it is safe to assume adiabatic propagation in our work.